OpenStudy (anonymous):

What is the equation of the oblique asymptote? h(x)=(x²-3x-4)/(x+1)

OpenStudy (zarkon):

x²-3x-4 factors

OpenStudy (dumbcow):

Factor x^2 -3x -4 = (x+1)(x-4)

OpenStudy (anonymous):

opps I meant to add a. y=x b. y=x+4 c. y=x-4 d. y=x²-3

OpenStudy (anonymous):

???

OpenStudy (zarkon):

you, pretty much, have been given the answer

OpenStudy (anonymous):

y=x-4

OpenStudy (anonymous):

you divide numerator by denominator. The result will be oblique asymptote

Directrix (directrix):

More information on oblique or slant asymptotes: http://www.purplemath.com/modules/asymtote3.htm http://www.purplemath.com/modules/asymnote.htm